BluStar Company has two service departments, Administration and Accounting, and two operating departments, Domestic and International. Administration costs are allocated on the basis of employees, and Accounting costs are allocated on the basis of number of transactions. A summary of BluStar operations follows.

Administration Accounting Domestic International
Employees 28 46 26
Transactions 32,000 25,000 100,000
Department direct costs $ 363,000 $ 145,000 $ 935,000 $ 3,590,000


Required:

a. Allocate the cost of the service departments to the operating departments using the direct method.
b. Allocate the cost of the service departments to the operating departments using the step method. Start with Administration.
c. Allocate the cost of the service departments to the operating departments using the reciprocal method

  • Required A
  • Required B
  • Required C

Allocate the cost of the service departments to the operating departments using the direct method. (Do not round intermediate calculations. Negative amounts should be indicated by a minus sign.)

To
From Administration Accounting Domestic International
Department costs $363,000 $145,000 $935,000 $3,590,000
Administration allocation (363,000)
Accounting allocation 0
Total cost $0 $145,000 $935,000 $3,590,000
    Answer #1

    - TO W Allocation of Cost using Direct Method To 7 From Administration Accounting Domestic International Total 5 No of Employ

    А Allocation of Cost using Step Method To - From Administration Accounting Domestic International Total - No of Employees 28

    А B U Allocation of Cost using Reciprocal Method Administration Accounting No of Employees % total Transactions % total 32000

    40 Accounting Cost = 145000+ 28% of (363000 + Cost from Accounting allocated to Admin) 41 Seeing the above table, we can obse
    Answer #2

    part A) Domestic 

    46+26= 72

    administration allocation: 363,000*46/72= 231,916.66666667 or 231.917

    25,000+100,000= 125,000

    accounting allocation: 145,000*25,000/125,000= 29,000


    international

    administration allocation:363,000- 231,917= 131,083

    accounting allocation: 145,000- 29,000= 116,000
    Answer #3

    part b) accounting 28+46+26= 100

    admin allocation 363,000*28/100= 101,640

    accounting allocation 145,000+ 101,640= 246,640 change to -246,640


    dometsic

    admin allocation 363,000*46/100= 166,980

    accounting allocation  246,640*25,000/125000= 49,328


    international

    admin allocation= 363,000- 101,640- 166,980= 94,380

    accounting allocation 246,640- 49,328= 197,312